Uno de los componentes principales del aceite de hierba de limón es un
compuesto con fórmula molecular C10H16. Cuando este compuesto se trata con
exceso de hidrógeno y un catalizador de platino, el producto es un alcano de fórmula
C10H20. ¿Qué puede concluir sobre la estructura de C10H16? Explica tu respuesta.

what information need?
Answer #1

Given main components of lemongrass oil is a compound with molecular formula C10H16

Now if we calculate the degree of unsaturation or Hydrogen defficiency i.e the total number of double bond and rings in the molecule,

then Degree of unsaturation = C + 1 - H/2 where C = number of C- atoms, H = number of H- atoms

Thus here in C10H16, Degree of unsaturation = C + 1 - H/2

= 10 + 1 - 16/2

= 11 - 8

= 3

It indicates that there is a possibility total C=C double bonds + ring = 3

Again when treated with excess hydrogen and a platinum catalyst, the product is an alkane of formula
C10H20

Now in this formula, Degree of unsaturation = C + 1 - H/2

= 10 + 1 - 20/2

   = 11 - 10

= 1

Since the given C10H20 is an alkane, thus there can't be any double bond here. Hence the Degree of unsaturation of Hydrogen deficiency is due to only 1 ring

Thus the possible structure of the compound C10H16 contain total two C=C double bonds + one ring.